Hello, can you please help me answer this :)
A father is four times his son's age now.
Five years ago he was seven times his son's age.
What is the father's age and son's age now.

Answer by sachi(548)

A father is four times his son's age now.
so if son is x yrs than father is 4x yrs
Five years ago he was seven times his son's age.
son was x-5 & father was 4x-5 yrs
so 4x-5=7(x-5)=7x-35
or 3x=30
or x=10
the father's age is 40 and son's age is 10 yrsnow.
ans